Calculate the mass (in grams) of 10 trillion silver atoms


1
Expert's answer
2021-09-23T02:37:05-0400

Atomic mass of silver (Ag)=107.87g

Avogadro mass of silver"=\frac{107.87}{1mole}\times \frac{1mole}{6.02\times 10^{23}atoms}=1.79\times10^{-22}g/atom"

"10trillion=10\times 10^{12}"

mass of 10 trillion of Ag atoms is"=\frac{1.79\times 10^{-22}g}{1.0Ag\space atom}\times (10\times 10^{12})Ag \space atoms=1.79\times 10^{-9}g"